What is the y-intercept for the exponential function y = bx ?
1
Expert's answer
2014-10-23T08:33:31-0400
Dear Lanaye. Are you sure that your question is correct? Function y=bx is not exponential, it is a linear function.
Maybe you meant y=ebx which is an exponential function. The y-intercept is a point where the graph of function crosses the y-axis, in other words, it is a point on the graph where x is zero. For y=bx, the y-intercept is (0,0), because y(0)=b*0=0. For y=ebx, the y-intercept is (0,1), because y(0)=eb*0=e0=1.